OpenStudy (anonymous):

There are 112 students on a team. they all ride the bus or bike to school. 4 buses were filled and 20 students ride their bikes. How many students were on each bus? What's the equation?

OpenStudy (anonymous):

Let the number of students on each bus be x. 4x + 20 = 112 4x = 112 - 20 = 92 x = 92/4 = 23 So, there were 23 students in each bus.

OpenStudy (anonymous):

Thanks alot

OpenStudy (anonymous):

can you give me a medal?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

how

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

Your welcome